The larger of two numbers is 1 more than twice the smaller. The sum of the numbers is 20 less than three times the larger. Find the numbers. I need this to be a system of two equations in two variables.
Answer by stanbon(75887)

The larger of two numbers is 1 more than twice the smaller. The sum of the numbers is 20 less than three times the larger. Find the numbers. I need this to be a system of two equations in two variables.
--------
Let x be the larger number; Let y be the smaller number.
EQUATIONS:
x-2y=1
x+y+20=3x
---------
Rewrite as
x-2y=1
2x-y=20
==========
Multiply the 1st equation by 2 to get:
2x-4y=2
Subtract this result from the 1st equation to get:
3y=18
y=6
Substitute that into the 1st equation to get:
x-2*6=1
x=13
-------
Solution: x=13; y=6
Cheers,
Stan H.